[Hallb-engineering] No Hall-B meeting today - here is an update
pasyuk at jlab.org
Mon May 30 11:28:03 EDT 2022
It is just a week left before the RGC run begins.
All shifts for the beginning of the run are covered. If you have not signed up for shifts yet, please do.
Reminder to those who have not been to JLab for a long time: register at least 7 days before arrival at JLab and make sure you have enough time to update your training.
The usual run documentation (COO, ESAD, RSAD, ERG) is posted on the run page: https://www.jlab.org/Hall-B/run-web/index_rgc.html
Everyone who will be taking shifts and/or working in the hall during RGC must read it. The yellow binder with paper copies of these documents and a signature list is in the counting house. When you first come to the counting house you must acknowledge that you have read and understood these documents by signing the signature list.
Familiarize yourself with CLAS12 subsystem manuals and procedures on the run wiki: https://wiki.jlab.org/clas12-run/index.php/Run_Group_C
To subsystem owners:
Please, review your system documentation and make changes if necessary.
Check your system and update its status in HCO. We would like to have most of the system checked and be in a "Ready" state by the end of this week.
Charge and check your on-call cell phone.
Feel free to contact me if you have any questions.
PDL for RGC,
From: Hallb-engineering <hallb-engineering-bounces at jlab.org> On Behalf Of Stepan Stepanyan
Sent: Monday, May 30, 2022 09:43
To: CLAS MEMBERS <clas_members at jlab.org>; hallb at jlab.org; Hallb-eng <hallb-engineering at jlab.org>
Cc: Achenbach, Prof. Dr. Patrick <achenbach at uni-mainz.de>
Subject: [Hallb-engineering] No Hall-B meeting today - here is an update
I see many are connected to the Hall-B meeting zoom. There is no meeting today. It is a holiday in the US, and the Lab is closed.
Here is a short update on the status of the hall and preparations for the run that will start in 10 days:
* target has been installed and aligned on the space frame. Cooldown has started
* no noise has been detected in CVT with all the pumps and the target RF running - great news!
* the second pair of the raster magnets are installed on the cart. Connections and the final testing will happen this week
* after a few troubles with the new power supply for the upstream Moller quad, the decision was made to reinstall and use the old one. Supplies for both quads have been tested, polarities checked, and are ready for the run.
* one of the requested additional LTCCs was installed in Sector 6. The second one is in the hall and will be installed in Sector 2 tomorrow.
* beam dump was open to check the beam blocker limit switched. Both switches have to be replaced engineering group, together with Rad. Cont. are getting ready for repair this week.
* other repair and maintenance work is progressing well on detectors (ECal/FTOF/CTOF/DC)
The biggest news for this week, I think, is the completion of the RICH-II assembly. Congratulations to the RICH team! The module is now getting ready to be mounted on the A-frame on Tuesday for transportation in the hall on Wednesday. Installation of the detector is scheduled for Thursday.
* RG-C: organizing schedule for online calibration
* RG-F: Request made for next pass-1 review - waiting for scheduling.
* RG-M: Additional calibration of 3 runs for FTOF getting underway based on recent pass-0 cooking. Work in progress on CVT alignment - alignment constants determined and testing in progress.
* DC: Systematic studies of pressure-dependence of time-to-distance calibration in progress.
Cooking status: RG-M: Cooking to support calibration requests; cooking for CVT alignment tests.
* DC: large alignment progress, see last week's software meeting: https://clasweb.jlab.org/wiki/index.php/Agendas,_minutes,_notes_and_related_documents
* CVT: efficiency improvements are in progress
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Hallb-engineering